OpenStudy (lilsis76):

Assume that inflation will increase the prices of everything at a rate of 5 percent a year over the next 10 years. In 3 years, what will be the price of a stereo that now costs $100? Round ur answer to the nearest hundreth.

OpenStudy (ybarrap):

Let P be the price at the end of the period of inflation. We assume that each year, inflation will increase by a factor of 5% / 10 years or .05/10: $$ P=P_0(1+.05/10)^n $$ Where \(P_0=$100\), is the price at the beginning (i.e \(n=0\)) and \(n\) is the number of years that have elapsed, 0,1,2,... What you want is P, the price at n=3, when the 3rd year hits. That's it!

OpenStudy (lilsis76):

what?! HAHA okay hold on @ybarrap haha I forgot to put multiple choice answers haha sorry these are my choices

OpenStudy (lilsis76):

a) $105 b) 115 c)115.52 d)115.76 e)150.

OpenStudy (lilsis76):

@ybarrap

OpenStudy (anonymous):

A = P (1+r)^3 but r = 0.05 not 0.05/10 A = 115.76

OpenStudy (ybarrap):

@douglaswinslowcooper is correct, I read as 5% over 10 years.

OpenStudy (lilsis76):

YAY!!! thank you lol

OpenStudy (anonymous):

Thank you both.

OpenStudy (ybarrap):

yw